### v1

# The Q-AMOEBA (CF) Polarizable Potential

Abstract
We present Q-AMOEBA (CF), an enhanced version of the Q-AMOEBA polarizable model that integrates a geometry-dependent charge flux (CF) term while designed for an explicit treatment of nuclear quantum effects (NQE). The inclusion of CF effects allows matching experimental data for the molecular structure of water in both gas and liquid phases, addressing limitations faced by most force fields. We show that Q-AMOEBA (CF) provides highly accurate results for a wide range of thermodynamical properties of liquid water. Using the computational efficiency of the adaptive Quantum Thermal Bath method, which accounts for NQE at a cost comparable to classical molecular dynamics, we evaluate the robustness and transferability of Q-AMOEBA (CF) by calculating hydration free energies of various ions and organic molecules. Finally, we apply this methodology to the alanine dipeptide and compute the corresponding dihedral angle potential of mean force and hydration free energy. Unexpectedly, the latter quantity displays significant NQE. These results pave the way to a finer understanding of their role in biochemical systems.
keywords:
American Chemical Society, LaTeXDepartment of Chemistry, University of Pittsburgh, Pittsburgh, Pennsylvania 15260, United States
1 Introduction
The pursuit of a realistic and accurate description of water across its diverse phases has long been central to computational molecular science, particularly in Monte Carlo (MC) and Molecular Dynamics (MD) simulations. Since the pioneering efforts of Stillinger and Rahman 1, 2, numerous computational water models have been developed 3, 4, 5, 6, each aiming to capture water’s intricate properties, such as its permanent dipole moment, non-monotonic density profile, and temperature-dependent dielectric constant to only name a few 7, 8. Early models, such as AMBER 9, CHARMM 10, OPLS 11, SPC 12, 13, and TIP4P 14, 15, 16, employed simplified representations, treating atoms as fixed point charges connected by harmonic bonds. While these simple force fields (FF) successfully capture a variety of molecular behaviors, they struggle to accurately represent complex phenomena like polarization and charge transfer, limiting their transferability to more complex environments. To overcome these limitations, polarizable FF17, 18, 19 such as TTMx-F 20, 21, AMOEBA 22, 23, 24, AMOEBA+ 25, 26, SIBFA 27, GEM* 28, HIPPO 29, MASTIFF 30, 31, MB-pol 32, 33, q-AQUA-pol 34, CMM 35, and others 36, were developed, introducing atomic polarizability to account for many-body effects. This enhancement allowed for more accurate simulations, particularly in systems characterized by strong electrostatic interactions, such as proteins and ionic liquids. However, these models are primarily parametrized for classical MD, and generally overlook the complexities introduced by Nuclear Quantum Effects (NQE), that can prove significant in systems involving light elements like hydrogen 37, 38, 39. Indeed, the omission of NQE, especially zero-point energy contributions, underestimates molecular vibrations, which impacts thermodynamic and structural properties and can lead to important discrepancies with experimental results 40, 41. In biological systems, where hydrogen plays a crucial role, addressing NQE is even more critical 42, 43, 44. NQE are often considered to be included implicitly in the potential energy surface (PES) of the most common FF, as their parameter values are fitted to reproduce experimental observables when performing MD simulations with classical nuclei. In this case, explicit incorporation of NQE requires a prior re-parameterization of the FF 45 to avoid double-counting of the quantum effects. NQE can be introduced explicitly via Path Integral Molecular Dynamics (PIMD): based on Feynman’s path integral formulation of quantum mechanics 46, PIMD maps quantum particles to a cyclic chain of classical particles (often called beads), making the method significantly more computationally demanding than classical MD. Consequently, PIMD has seen limited use, although several water potentials have been developed using this method 47, 48, 49, 50, 51, 52. Different approaches have been proposed to lower PIMD’s computational overhead by reducing the required number of beads53, 54, 55, 56, 57, 58, but the cost generally remains high compared to classical MD. As an alternative to PIMD, the Quantum Thermal Bath (QTB) method 59, 60 approximates the quantum statistical distribution with a computational cost comparable to classical (MD). Though the original approach suffers from zero-point energy leakage, limiting its accuracy, a solution was found with the adaptive QTB (adQTB), that enforces the quantum fluctuation-dissipation theorem in a systematic way61. This method has proven effective in computing NQE in liquid water 62 and developing a refined model based on the polarizable AMOEBA PES, named Q-AMOEBA 63. While Q-AMOEBA was shown to provide a remarkably accurate description of various properties of liquid water63, its charge distribution does not fully respond to local geometry changes, as it does not account for intramolecular charge transfer or Charge Flux (CF) 64, 65, 26, essential for accurate water modeling 66, 67. Furthermore, previous polarizable models that include a CF term25 were not designed to explicitly account for NQE.
Therefore, in this paper, we introduce a geometry-dependent CF term into our previously developed Q-AMOEBA model. The new model is referred to as Q-AMOEBA (CF) and was implemented within the GPU-accelerated Tinker-HP software 68, 69 and leverages its Quantum-HP NQE scalable package 70. This enhancement seeks to improve the accuracy and transferability of Q-AMOEBA across diverse environments (gas phase, bulk systems, and interfaces). We will outline our parametrization procedure, and compare structural and thermodynamic properties of liquid water with the original AMOEBA-03 model 22, before exploring the hydration free energies (HFE) of different ions and small molecules. Finally, to further demonstrate the robustness of the new Q-AMOEBA (CF) model, we turn to a larger compound, the alanine dipeptide, for which we compute the HFE and the dihedral angle potential of mean force, taking advantage of the efficient enhanced sampling methods implemented in Tinker-HP.
2 Parametrization procedure
The AMOEBA potential energy surface is composed of bonded and nonbonded energy terms, detailed in refs 22, 24, 25. Compared to the previous Q-AMOEBA model 63 (and to the original AMOEBA-03), different parameters were modified. First, the inclusion of the Charge Flux (CF) correction, adapted from the AMOEBA+ model 26, which reproduces intramolecular charge reorganization with geometry deformations. In particular, this correction was shown to improve the agreement with the molecular bond angle experimental values across the gas and liquid phases, thereby addressing a common issue in water force fields 22, 25, 26. The CF parameters were initialized from the AMOEBA+ parameters 26 and further adjusted to recover the experimental bond length and angle in simulations with explicit treatment of NQE. Then, the bond stretching force parameter was slightly reduced to improve the agreement of the simulated infrared (IR) absorption spectra with experimental data. Finally, the van der Waals (vdW) parameters were fine-tuned using as target property the experimental density at four temperatures (249.15 K, 277.15 K, 298.15 K, and 341.15 K). The final list of parameters is provided in supporting information (SI). As discussed below, the incorporation of the CF term and the subsequent adjustments enable the use of a consistent set of intramolecular parameters across gas and liquid phases, with a good match to the experimental data for molecular structure properties (see Table 1). In addition, since adQTB and PIMD simulations lead to slightly different values for the liquid water density62, 63, two distinct vdW parameter sets were developed for each of the two methods. Unless stated otherwise, the results provided here for the Q-AMOEBA (CF) model are obtained from adQTB simulations (using the adQTB parameters set). The PIMD results are generally very similar, as shown from the detailed analysis available in SI. The SI also provides an extensive comparison with an updated version of our previous Q-AMOEBA model (with reduced bond stretching force parameter but without CF term).
Liquid water simulations used 4000 molecules in a cubic box with periodic boundary conditions. The vdW cutoff was 12 Å, electrostatic interactions were computed with the Smooth Particle Mesh Ewald (SPME) method 71, 72 using a 7 Å real-space cutoff and a 60 × 60 × 60 grid. A BAOAB-RESPA integrator 73 with 0.2 fs and 2 fs time steps for bonded and nonbonded interactions, respectively, was used. PIMD simulations employed 32 beads with the Thermostated Ring Polymer Molecular Dynamics (TRPMD) method with a mild Langevin thermostat on the centroid with a friction coefficient of ps-1. In contrast, the adQTB used a higher friction coefficient of 20 ps-1.
Gas Phase | Liquid Phase | |||
(Å) | (deg) | (Å) | (deg) | |
Q-AMOEBA (CF) | 0.957 | 104.50 | 0.983 | 105.15 |
AMOEBA-03 | 0.957 | 108.50 | 0.968 | 106.10 |
Exp. | 0.957 | 104.52 | 0.97 | 105.1 |
3 Results and Discussion
3.1 Binding energies of Water Clusters
Table 2 shows the Binding Energies (BE) for various water clusters, comparing the Q-AMOEBA (CF) model with AMOEBA-03 and ab initio methods 74, 76, 77.
This allows us to validate the new parameters against the original AMOEBA-03 and ab initio BE. The inclusion of CF has reduced the Root Mean Squared Error (RMSE) to 0.37, compared to 0.81 for AMOEBA-03.
Table IV in SI demonstrates that Q-AMOEBA (CF) reduces the RMSE by about 25% compared to AMOEBA-03 for larger water clusters, regardless of the method (PIMD or adQTB) used for including NQEs. This highlights the importance of precise monomer geometry in a flexible water model for faithfully reproducing the intricate PES of water clusters.
3.2 Liquid properties of water
In bulk liquid water, the H2O molecules’ structural dynamics generate a globally balanced and symmetric environment, resulting in most molecules having an average net charge close to zero. Therefore, Q-AMOEBA (CF) only considers intramolecular charge transfers, which are described by the Charge Flux term. This CF term couples to the many-body polarization of the system and allows to capture the variations of the molecular dipole with its chemical environment. As a result of this improved description of polarization effects, Q-AMOEBA (CF) predicts a change in dipole moment values from 1.84 D in the gas phase to 2.77 D in the liquid phase. These results align closely with both ab initio MD calculations 78 and experimental observations 79, 80, which report a gas-phase dipole moment of 1.85 D and a liquid-phase near 2.9 D. In contrast, AMOEBA-03 predicts a similar value for the liquid phase (2.75 D) but underestimates the gas-phase dipole (1.77 D).
More generally, Table 1 presents the properties of intramolecular water structures and shows that the Q-AMOEBA (CF), with its geometry-dependent term, aligns with experimental data across gas and liquid phases.
NQE also significantly influence vibrational spectroscopy, affecting peak positions and intensities in IR spectra. Simulations not accounting for these effects often deviate from experimental results. Figure 1-(a) shows IR spectra for Q-AMOEBA (CF) with a deconvolution procedure applied to correct for broad spectral lines due to high friction coefficients used in the dynamics, as previously reported 62. The blue shift in intramolecular peak positions between AMOEBA-03 and Q-AMOEBA (CF) is a well known impact of NQE 39. The PIMD curves (SI-Fig.4) display only marginal differences, suggesting that the observed variations in this region stem from the enhanced PES rather than the approach employed for computing NQE. Using a lower gas-phase bond stretching force in our updated FF leads to peak positions that closely match experimental data without affecting liquid thermodynamic properties. Additionally, CF induces red shifts in the OH stretching region, aligning with experimental observations 81, 82, 83. This contrasts with AMOEBA+ 25, which shows a blue shift, likely due to the absence of NQE.
Thermodynamic properties such as density, heat capacity, thermal expansion, and isothermal compressibility were assessed from 249.15 K to 369.15 K at 1 atm (see Figure 1-(b)-(g)). Q-AMOEBA (CF) accurately reproduces the unusual bell-shaped density curve of water, while AMOEBA-03 shows a notable decrease in density at low temperatures, deviating from experimental observations. NQE influence water’s density maximum due to oxygen’s electronegativity, creating polarized OH bonds and directional hydrogen bonding that stabilize ice’s tetrahedral structure. The observed density maximum shifts for different isotopes (277.15 K for H2O, 284.15 for D2O, and 286.15 for T2O) 84 indicate an enhanced hydrogen bond network with increasing isotope mass. Notably, the inclusion of NQE and CF brings the predicted maximum density closer to the experimental value, with Q-AMOEBA (CF) at 279.5 K compared to AMOEBA-03’s higher value of 287 K.
The enthalpy of vaporization (Hvap) is crucial for quantifying hydrogen bonds and NQE in water. Q-AMOEBA (CF) achieves chemical accuracy for Hvap across all temperatures, surpassing AMOEBA-03 and previous Q-AMOEBA versions 63. This improvement, evident in the correct slope, underscores NQE’s essential role. Notably, H was not used in parameter fitting. Despite this, Q-AMOEBA (CF) accurately captures thermodynamic observables strongly influenced by NQE, demonstrating its robustness even with fewer fitted bulk properties. This underscores the necessity of incorporating both NQE and CF to accurately describe water’s thermodynamic properties.
While CF’s impact on the dielectric constant is subtle (Figure 1-(g)), it significantly improves the description of molecular charge distribution. By enhancing the accuracy of dipole moment calculations, Q-AMOEBA (CF) demonstrates superior capability in modeling water’s local geometry and polarization effects, making it a crucial advancement in understanding water’s complex behavior.
Finally, Q-AMOEBA (CF) predicts a self-diffusion coefficient of 0.80 ± 0.01 × 105 cm2/s at 298.15 K under high friction coefficients, which is lower than experimental values. Adjusting to a lower friction coefficient with thermostatted ring-polymer MD (TRPMD) yields values closer to experiments (2.42 ± 0.01 × 105 cm2/s). In comparison, AMOEBA’s value of 2.02 × 105 cm2/s demonstrates that including NQE accelerates diffusion compared to classical dynamics. Further studies on ice I and heavy water in SI expand on the role of NQE with Q-AMOEBA (CF).
3.3 Hydration Free Energy
H2O | F- | Na+ | Mg2+ | Cl- | K+ | Ca2+ | Br- | I- | |
Q-AMOEBA (CF) | -5.79 (-5.69) | -116.1 | -88.5 | -427.7 | -85.7 | -71.4 | -351.2 | -79.6 | -70.8 |
AMOEBA-03 | - 5.78 | -114.7 | -89.7 | -432.4 | -84.1 | -72.3 | -354.6 | -77.7 | -69.3 |
Expt. | -6.32 | -119.7 | -88.7 | -435 | -89.1 | -71.2 | -356.8 | -82.7 | -74.3 |
The lower transferability of models are usually compensated by employing different parameters for different situations such as proteins or on how the long-range electrostatic interactions is computed. Table 3 shows the different Hydration Free Energy (HFE) computed with Q-AMOEBA (CF) and AMOEBA-03 compared to experimental results of different ions and water. The HFE were computed using a free energy perturbation procedure based on twenty-one thermodynamic states in the NVT ensemble at 298.15 K. The detailed MD window parameter can be found in SI. Each thermodynamical state was thermalized for 1.5 ns and run for 5 ns. The Bennett Acceptance ratio (BAR) method 91 and the Path Integral BAR (PI-BAR) for PIMD results 70 was then used to estimate the free energy. The final free energy was taken as the sum over all windows were 1.9 kcal.mol-1 was added for each ions 90. The original AMOEBA-03 yields similar result than Q-AMOEBA (CF) for water as one may see in Table 3. However, AMOEBA-03 was fitted to implicitly include NQE. Therefore, obtaining values as close as the original AMOEBA-03 indicated that some errors in the PES have been removed, resulting in an overall improved PES with Q-AMOEBA (CF). This discrepancy between the experimental value and Q-AMOEBA (CF) is expected due to the underestimation of the enthalpy (see Figure 1-(c)) and the strength of H bonds. Looking at Table 3, one can notice an overall improvement in the HFE values for the different ions, with the exception of Mg2+ and Ca2+, compared to the original AMOEBA-03.
To further assess the transferability of Q-AMOEBA (CF), we also proceed to compute the HFE of approximately 40 small organic compounds which were parametrized using Poltype2 92. Figure 2 displays the scatter plot of Q-AMOEBA (CF) models with AMOEBA-03 against experimental values alongside their respective RMSE. The original AMOEBA-03 model exhibits better agreement with experimental data (0.58 for AMOEBA-03 and 0.94 for Q-AMOEBA (CF) model). Despite this, Q-AMOEBA (CF) model has a strong correlation coefficient (0.97), albeit with slightly higher RMSE values. Considering the explicit inclusion of NQE in Q-AMOEBA (CF), the higher RMSE values are reasonable, as these effects introduce additional complexity to the model.
3.4 The Alanine Dipeptide
A longstanding question in the field of biomolecular simulation concerns the quantitative and qualitative impacts of NQE 93 in proteins and in protein-ligand systems. Such calculations present inherent challenges due to the high number of conformational states accessible. In this context, we focus on a prototypical system, the alanine dipeptide, a relatively simple yet insightful model that captures key sampling challenges also present in larger proteins. Indeed, it embodies sampling difficulties typical of larger protein with its well known metastable states characterized by the two dihedral angles and (shown in Figure 3) serves as an ideal model for exploring NQE impacts.
To overcome these, we first resort to the Adaptive Biasing force technique 94, 95 using the Colvars library 96 with both the and as collective variables in order to recover the 2 dimensional free energy surface as a function of (,) with NQE using adQTB and without NQE. Details of the simulation, and input files are provided in SI. Our simulations (available in SI) reveal only marginal differences in the free energy surfaces, suggesting that NQE have a minimal effect in this particular system. This highlights the robustness of both the adQTB method and our model in handling complex biological systems. We then investigate the impact of NQE on the hydration free energy of alanine dipeptide using the Q-AMOEBA (CF) model with NQE and its classical counterpart. The results, shown in Table 4, are compared with those from the AMOEBA-03 force field. To compute this, we leverage the recently introduced Lambda-ABF technique within the Tinker-HP framework, coupled to Colvars doi:10.1021/acs.jctc.3c01249. This framework allows for efficient sampling of the alchemical variable while biasing other slow degrees of freedom, such as and . Given that is known to be associated with the largest free energy barrier, we perform 2D ABF simulations with respect to both and , yielding 2D free energy surfaces. The hydration free energy of alanine dipeptide is then recovered by marginalizing with respect to . The hydration free energy computed with the Q-AMOEBA (CF) model and NQE is approximately 4 kcal/mol lower than that obtained using the classical version of the model, indicating that NQE reduces the hydrophilic character of the alanine dipeptide. While electrostatic contributions remain consistent across all models, vdW interactions vary significantly due to the effects of NQE (Figure available in SI). This variation stems from quantum-induced atomic delocalization, particularly in light atoms such as hydrogen, which broadens the range of accessible molecular configurations and enhances entropic contributions. As a result, the strength of vdW interactions changes, while electrostatic interactions—primarily determined by fixed charge distributions and distances—are less influenced. Ultimately, the enhanced sampling enabled by NQE leads to greater entropic contributions to the hydration free energy, causing discrepancies between the quantum and classical models. Interestingly, the AMOEBA-03 results closely matched those of the classical Q-AMOEBA (CF) model, indicating that neglecting NQE, especially in systems where vdW forces play a significant role, can lead to inaccurate free energy estimates.
Hydration Free Energy | |
Q-AMOEBA (CF) - adQTB | -8.37 |
Q-AMOEBA (CF) - Classical | -13.50 |
AMOEBA-03 | -12.25 |
4 Conclusion
In this work, we introduced Q-AMOEBA (CF), an enhanced version of the Q-AMOEBA model that incorporates a geometry-dependent CF term and explicitly accounts for NQE via the adQTB method. The CF correction addresses key limitations of traditional FF by refining molecular geometries to align bond lengths and angles with experimental data across both gas and liquid phases. Q-AMOEBA (CF) is designed to be used in simulations with explicit inclusion of NQE and provides a step forward towards a more accurate description of molecular properties, particularly in the context of water and biological systems. This enables the prediction of accurate hydration free energies of organic molecules and the study of more complex systems, such as the alanine dipeptide, overall with improved efficiency and precision, showcasing the broad applicability and versatility of Q-AMOEBA (CF). While other models include NQE, their computational expense often limits their use in large systems. The Q-AMOEBA (CF) model offers a balance between accuracy and computational efficiency, marking a milestone in the development of next-generation polarizable force fields for simulating biological and biochemical systems. It is also important to note that the vdW interactions in biological systems may require further attention in order to explicitly use NQE at large scale. This parametrization task will be the subject of further research in link with the biological importance of NQE in biology and will be effectively addressed using machine learning techniques, which hold great potential in training models that adapt the vdW parameters for biological contexts97.
Acknowledgments
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 810367), project EMC2 (JPP). Computations have been performed at IDRIS (Jean Zay) on GENCI Grants: no A0150712052 (J.-P. P.).

### v2

@@
 Different approaches have been proposed to lower PIMD’s computational overhead by reducing the required number of beads53, 54, 55, 56, 57, 58, but the cost generally remains high compared to classical MD.
-As an alternative to PIMD, the Quantum Thermal Bath (QTB) method 59, 60 approximates the quantum statistical distribution with a computational cost comparable to classical (MD).
+As an alternative to PIMD, the Quantum Thermal Bath (QTB) method 59, 60 approximates the quantum statistical distribution with a computational cost comparable to classical MD.
 Though the original approach suffers from zero-point energy leakage, limiting its accuracy, a solution was found with the adaptive QTB (adQTB), that enforces the quantum fluctuation-dissipation theorem in a systematic way61.
@@
 The inclusion of CF has reduced the Root Mean Squared Error (RMSE) to 0.37, compared to 0.81 for AMOEBA-03.
-Table IV in SI demonstrates that Q-AMOEBA (CF) reduces the RMSE by about 25% compared to AMOEBA-03 for larger water clusters, regardless of the method (PIMD or adQTB) used for including NQEs.
+Table IV in SI demonstrates that Q-AMOEBA (CF) reduces the RMSE by about 25% compared to AMOEBA-03 for larger water clusters, regardless of the method (PIMD or adQTB) used for including NQEs during the parametrization.
 This highlights the importance of precise monomer geometry in a flexible water model for faithfully reproducing the intricate PES of water clusters.
@@
 Simulations not accounting for these effects often deviate from experimental results.
-Figure 1-(a) shows IR spectra for Q-AMOEBA (CF) with a deconvolution procedure applied to correct for broad spectral lines due to high friction coefficients used in the dynamics, as previously reported 62.
-The blue shift in intramolecular peak positions between AMOEBA-03 and Q-AMOEBA (CF) is a well known impact of NQE 39.
-The PIMD curves (SI-Fig.4) display only marginal differences, suggesting that the observed variations in this region stem from the enhanced PES rather than the approach employed for computing NQE.
-Using a lower gas-phase bond stretching force in our updated FF leads to peak positions that closely match experimental data without affecting liquid thermodynamic properties.
-Additionally, CF induces red shifts in the OH stretching region, aligning with experimental observations 81, 82, 83.
-This contrasts with AMOEBA+ 25, which shows a blue shift, likely due to the absence of NQE.
+Figure 1-(a) shows IR spectra for Q-AMOEBA (CF) with a deconvolution procedure applied to correct for the spectral broadening caused by the high friction coefficient used in adQTB simulations, as previously reported 62.
+The explicit inclusion of NQE generally tends to red-shift the intramolecular peak frequencies39, which is consistent with the results of Fig.
+1-(a), comparing Q-AMOEBA (CF) and AMOEBA-03.
+This trend is independent on the method used for the inclusion of NQE and the PIMD curves provided in SI-Fig.4 only display marginal discrepancies with respect to the adQTB results of Fig.
+1-(a).
+Using a lower gas-phase bond stretching force in our updated FF leads to a stretching peak position (at 3400 cm-1) that closely match experimental data, without affecting liquid thermodynamic properties.
+Additionally, the CF term induces a notable red shift and intensity reduction of the angle bending peak (at 1600 cm-1) , aligning with the experimental spectrum 81, 82, 83, and in agreement with previous observations for the AMOEBA+ 25 force field.
 Thermodynamic properties such as density, heat capacity, thermal expansion, and isothermal compressibility were assessed from 249.15 K to 369.15 K at 1 atm (see Figure 1-(b)-(g)).
 Q-AMOEBA (CF) accurately reproduces the unusual bell-shaped density curve of water, while AMOEBA-03 shows a notable decrease in density at low temperatures, deviating from experimental observations.
-NQE influence water’s density maximum due to oxygen’s electronegativity, creating polarized OH bonds and directional hydrogen bonding that stabilize ice’s tetrahedral structure.
-The observed density maximum shifts for different isotopes (277.15 K for H2O, 284.15 for D2O, and 286.15 for T2O) 84 indicate an enhanced hydrogen bond network with increasing isotope mass.
+NQE influence water’s density maximum due to oxygen’s electronegativity, which creates polarized OH bonds and directional hydrogen bonding that stabilizes local tetrahedral structures (similar to that found in ice).
+The observed density maximum shifts for different isotopes (277.13K for H2O, 284.34 K for D2O, and 286.55 K for T2O) 84, 39 indicate an enhanced hydrogen bond network with increasing isotope mass.
 Notably, the inclusion of NQE and CF brings the predicted maximum density closer to the experimental value, with Q-AMOEBA (CF) at 279.5 K compared to AMOEBA-03’s higher value of 287 K.
-The enthalpy of vaporization (Hvap) is crucial for quantifying hydrogen bonds and NQE in water.
+The enthalpy of vaporization (Hvap) is an other characteristic quantity that is known to be strongly affected by NQE in water.
 Q-AMOEBA (CF) achieves chemical accuracy for Hvap across all temperatures, surpassing AMOEBA-03 and previous Q-AMOEBA versions 63.
-This improvement, evident in the correct slope, underscores NQE’s essential role.
+This improvement, evident in the correct slope obtained with increasing temperature, underscores NQE’s essential role.
 Notably, H was not used in parameter fitting.
@@
 By enhancing the accuracy of dipole moment calculations, Q-AMOEBA (CF) demonstrates superior capability in modeling water’s local geometry and polarization effects, making it a crucial advancement in understanding water’s complex behavior.
-Finally, Q-AMOEBA (CF) predicts a self-diffusion coefficient of 0.80 ± 0.01 × 105 cm2/s at 298.15 K under high friction coefficients, which is lower than experimental values.
-Adjusting to a lower friction coefficient with thermostatted ring-polymer MD (TRPMD) yields values closer to experiments (2.42 ± 0.01 × 105 cm2/s).
-In comparison, AMOEBA’s value of 2.02 × 105 cm2/s demonstrates that including NQE accelerates diffusion compared to classical dynamics.
-Further studies on ice I and heavy water in SI expand on the role of NQE with Q-AMOEBA (CF).
+Finally, Q-AMOEBA (CF) predicts a self-diffusion coefficient of cm2/s at 298.15 K, in good agreement with experiments.
+In comparison, AMOEBA-03’s value of 2.02 × 105 cm2/s demonstrates that including NQE accelerates diffusion compared to classical dynamics.
+Note that the diffusion coefficient was calculated for Q-AMOEBA (CF) using thermostatted ring-polymer MD (TRPMD) as the adQTB method was shown in previous studies62, 63 to underestimate diffusion due to the high friction coefficients required to control zero-point energy leakage (the adQTB value for the coefficient is indeed 0.80 ± 0.01 × 105 cm2/s).
+Further studies on ice I and heavy water provided in SI expand on the role of NQE with Q-AMOEBA (CF).
+. Corresponding radial distribution functions are available in SI.
 3.3 Hydration Free Energy
 | H2O | F- | Na+ | Mg2+ | Cl- | K+ | Ca2+ | Br- | I- | |
-Q-AMOEBA (CF) | -5.79 (-5.69) | -116.1 | -88.5 | -427.7 | -85.7 | -71.4 | -351.2 | -79.6 | -70.8 |
+| Q-AMOEBA (CF) | -5.79 (-5.69) | -116.1 | -88.5 | -434.4 | -85.7 | -71.4 | -355.0 | -79.6 | -70.8 |
 | AMOEBA-03 | - 5.78 | -114.7 | -89.7 | -432.4 | -84.1 | -72.3 | -354.6 | -77.7 | -69.3 |
 | Expt. | -6.32 | -119.7 | -88.7 | -435 | -89.1 | -71.2 | -356.8 | -82.7 | -74.3 |
-The lower transferability of models are usually compensated by employing different parameters for different situations such as proteins or on how the long-range electrostatic interactions is computed.
-Table 3 shows the different Hydration Free Energy (HFE) computed with Q-AMOEBA (CF) and AMOEBA-03 compared to experimental results of different ions and water.
+Table 3 shows the Hydration Free Energies (HFE) computed with Q-AMOEBA (CF) and AMOEBA-03 compared to experimental results for different ions and for water.
 The HFE were computed using a free energy perturbation procedure based on twenty-one thermodynamic states in the NVT ensemble at 298.15 K.
@@
 The Bennett Acceptance ratio (BAR) method 91 and the Path Integral BAR (PI-BAR) for PIMD results 70 was then used to estimate the free energy.
-The final free energy was taken as the sum over all windows were 1.9 kcal.mol-1 was added for each ions 90.
-The original AMOEBA-03 yields similar result than Q-AMOEBA (CF) for water as one may see in Table 3.
-However, AMOEBA-03 was fitted to implicitly include NQE.
-Therefore, obtaining values as close as the original AMOEBA-03 indicated that some errors in the PES have been removed, resulting in an overall improved PES with Q-AMOEBA (CF).
-This discrepancy between the experimental value and Q-AMOEBA (CF) is expected due to the underestimation of the enthalpy (see Figure 1-(c)) and the strength of H bonds.
-Looking at Table 3, one can notice an overall improvement in the HFE values for the different ions, with the exception of Mg2+ and Ca2+, compared to the original AMOEBA-03.
+The final free energy was taken as the sum over all windows, where 1.9 kcal.mol-1 was added for each ions 90.
+For water, the Q-AMOEBA (CF) model with explicit inclusion of NQE yields results that are very close to that of the original AMOEBA-03 (using classical MD), with a slight underestimation of the corresponding HFE with respect to experiment.
+Table 3, shows a clear overall improvement in the HFE values for the different ions.
+Note that for monovalent ions, this improvement was obtained with the original AMOEBA parameters90, while the Thole parameters have been slightly adjusted for Mg2+ and Ca2+ to correctly account for NQE with Q-AMOEBA (CF).
+The new parameters are available in SI.
 To further assess the transferability of Q-AMOEBA (CF), we also proceed to compute the HFE of approximately 40 small organic compounds which were parametrized using Poltype2 92.
@@
 The original AMOEBA-03 model exhibits better agreement with experimental data (0.58 for AMOEBA-03 and 0.94 for Q-AMOEBA (CF) model).
-Despite this, Q-AMOEBA (CF) model has a strong correlation coefficient (0.97), albeit with slightly higher RMSE values.
+Despite this, the Q-AMOEBA (CF) model has a strong correlation coefficient (0.97), albeit with slightly higher RMSE values.
 Considering the explicit inclusion of NQE in Q-AMOEBA (CF), the higher RMSE values are reasonable, as these effects introduce additional complexity to the model.
@@
 In this context, we focus on a prototypical system, the alanine dipeptide, a relatively simple yet insightful model that captures key sampling challenges also present in larger proteins.
-Indeed, it embodies sampling difficulties typical of larger protein with its well known metastable states characterized by the two dihedral angles and (shown in Figure 3) serves as an ideal model for exploring NQE impacts.
-To overcome these, we first resort to the Adaptive Biasing force technique 94, 95 using the Colvars library 96 with both the and as collective variables in order to recover the 2 dimensional free energy surface as a function of (,) with NQE using adQTB and without NQE.
+Indeed, it embodies sampling difficulties typical of larger proteins and its well known metastable states characterized by the two dihedral angles and (shown in Figure 3) serves as an ideal model for exploring NQE impacts.
+To overcome these, we first resort to the Adaptive Biasing force (ABF) technique 94, 95 using the Colvars library 96 with both the and as collective variables in order to recover the 2 dimensional free energy surface as a function of (,) with NQE using adQTB and without NQE.
 Details of the simulation, and input files are provided in SI.

### v3

@@
 American Chemical Society, LaTeXDepartment of Chemistry, University of Pittsburgh, Pittsburgh, Pennsylvania 15260, United States
-1 Introduction
 The pursuit of a realistic and accurate description of water across its diverse phases has long been central to computational molecular science, particularly in Monte Carlo (MC) and Molecular Dynamics (MD) simulations.
@@
 While these simple force fields (FF) successfully capture a variety of molecular behaviors, they struggle to accurately represent complex phenomena like polarization and charge transfer, limiting their transferability to more complex environments.
-To overcome these limitations, polarizable FF17, 18, 19 such as TTMx-F 20, 21, AMOEBA 22, 23, 24, AMOEBA+ 25, 26, SIBFA 27, GEM* 28, HIPPO 29, MASTIFF 30, 31, MB-pol 32, 33, q-AQUA-pol 34, CMM 35, and others 36, were developed, introducing atomic polarizability to account for many-body effects.
+To overcome these limitations, polarizable FF17, 18, 19 such as TTMx-F 20, 21, AMOEBA 22, 23, 24, AMOEBA+ 25, 26, SIBFA 27, GEM* 28, HIPPO 29, MASTIFF 30, 31, CMM 32, and others 33, were developed, introducing atomic polarizability to account for many-body effects.
 This enhancement allowed for more accurate simulations, particularly in systems characterized by strong electrostatic interactions, such as proteins and ionic liquids.
-However, these models are primarily parametrized for classical MD, and generally overlook the complexities introduced by Nuclear Quantum Effects (NQE), that can prove significant in systems involving light elements like hydrogen 37, 38, 39.
-Indeed, the omission of NQE, especially zero-point energy contributions, underestimates molecular vibrations, which impacts thermodynamic and structural properties and can lead to important discrepancies with experimental results 40, 41.
-In biological systems, where hydrogen plays a crucial role, addressing NQE is even more critical 42, 43, 44.
+However, these models are primarily parametrized for classical MD, and generally overlook the complexities introduced by Nuclear Quantum Effects (NQE), that can prove significant in systems involving light elements like hydrogen 34, 35, 36.
+Some recent force fields, such as MB-pol 37, 38, q-AQUA-pol 39, ARROW 40; and neural network potentials such as FENNIX 41 and FENNIX-Bio142 have addressed this limitation by incorporating NQE via various strategies.
+Indeed, the omission of NQE, especially zero-point energy contributions, underestimates molecular vibrations, which impacts thermodynamic and structural properties and can lead to important discrepancies with experimental results 43, 44.
+In biological systems, where hydrogen plays a crucial role, addressing NQE is even more critical 45, 46, 47.
 NQE are often considered to be included implicitly in the potential energy surface (PES) of the most common FF, as their parameter values are fitted to reproduce experimental observables when performing MD simulations with classical nuclei.
-In this case, explicit incorporation of NQE requires a prior re-parameterization of the FF 45 to avoid double-counting of the quantum effects.
-NQE can be introduced explicitly via Path Integral Molecular Dynamics (PIMD): based on Feynman’s path integral formulation of quantum mechanics 46, PIMD maps quantum particles to a cyclic chain of classical particles (often called beads), making the method significantly more computationally demanding than classical MD.
-Consequently, PIMD has seen limited use, although several water potentials have been developed using this method 47, 48, 49, 50, 51, 52.
-Different approaches have been proposed to lower PIMD’s computational overhead by reducing the required number of beads53, 54, 55, 56, 57, 58, but the cost generally remains high compared to classical MD.
-As an alternative to PIMD, the Quantum Thermal Bath (QTB) method 59, 60 approximates the quantum statistical distribution with a computational cost comparable to classical MD.
-Though the original approach suffers from zero-point energy leakage, limiting its accuracy, a solution was found with the adaptive QTB (adQTB), that enforces the quantum fluctuation-dissipation theorem in a systematic way61.
-This method has proven effective in computing NQE in liquid water 62 and developing a refined model based on the polarizable AMOEBA PES, named Q-AMOEBA 63.
-While Q-AMOEBA was shown to provide a remarkably accurate description of various properties of liquid water63, its charge distribution does not fully respond to local geometry changes, as it does not account for intramolecular charge transfer or Charge Flux (CF) 64, 65, 26, essential for accurate water modeling 66, 67.
-Furthermore, previous polarizable models that include a CF term25 were not designed to explicitly account for NQE.
+In this case, explicit incorporation of NQE requires a prior re-parameterization of the FF 48 to avoid double-counting of the quantum effects.
+NQE can be introduced explicitly via Path Integral Molecular Dynamics (PIMD): based on Feynman’s path integral formulation of quantum mechanics 49, PIMD maps quantum particles to a cyclic chain of classical particles (often called beads), making the method significantly more computationally demanding than classical MD.
+Consequently, PIMD has seen limited use, although several water potentials have been developed using this method 50, 51, 52, 53, 54, 55.
+Different approaches have been proposed to lower PIMD’s computational overhead by reducing the required number of beads56, 57, 58, 59, 60, 61, but for most applications, the cost generally remains significantly higher than that of classical MD simulations.
+As an alternative to PIMD, the Quantum Thermal Bath (QTB) method 62, 63 approximates the quantum statistical distribution with a computational cost comparable to classical MD.
+Though the original approach suffers from zero-point energy leakage, limiting its accuracy, a solution was found with the adaptive QTB (adQTB), that enforces the quantum fluctuation-dissipation theorem in a systematic way64.
+This method has proven effective in computing NQE in liquid water 65 and developing a refined model based on the polarizable AMOEBA PES, named Q-AMOEBA 66.
+While Q-AMOEBA was shown to provide a remarkably accurate description of various properties of liquid water66, its charge distribution does not fully respond to local geometry changes, as it does not account for intramolecular charge transfer or Charge Flux (CF) 67, 68, 26, essential for accurate water modeling 69, 70.
+Furthermore, previous polarizable models that include a CF term25 were not designed to explicitly account for NQEs.
 Therefore, in this paper, we introduce a geometry-dependent CF term into our previously developed Q-AMOEBA model.
-The new model is referred to as Q-AMOEBA (CF) and was implemented within the GPU-accelerated Tinker-HP software 68, 69 and leverages its Quantum-HP NQE scalable package 70.
+The new model is referred to as Q-AMOEBA (CF) and was implemented within the GPU-accelerated Tinker-HP software 71, 72 and leverages its Quantum-HP NQE scalable package 73.
 This enhancement seeks to improve the accuracy and transferability of Q-AMOEBA across diverse environments (gas phase, bulk systems, and interfaces).
@@
 Finally, to further demonstrate the robustness of the new Q-AMOEBA (CF) model, we turn to a larger compound, the alanine dipeptide, for which we compute the HFE and the dihedral angle potential of mean force, taking advantage of the efficient enhanced sampling methods implemented in Tinker-HP.
-2 Parametrization procedure
-The AMOEBA potential energy surface is composed of bonded and nonbonded energy terms, detailed in refs 22, 24, 25.
-Compared to the previous Q-AMOEBA model 63 (and to the original AMOEBA-03), different parameters were modified.
+The AMOEBA potential energy surface is composed of bonded and nonbonded energy term can be expressed as the sum of bonded and non-bonded energy terms 22, 24, 25:
+Compared to the previous Q-AMOEBA model 66 (and to the original AMOEBA-03), different parameters were modified.
 First, the inclusion of the Charge Flux (CF) correction, adapted from the AMOEBA+ model 26, which reproduces intramolecular charge reorganization with geometry deformations.
@@
 The CF parameters were initialized from the AMOEBA+ parameters 26 and further adjusted to recover the experimental bond length and angle in simulations with explicit treatment of NQE.
+Compared to AMOEBA+, only the angular CF parameter (jΘ) was significantly modified, increasing by an order of magnitude, from 0.0020 to 0.0433 e.degree, to account for the enhanced angular fluctuations induced by NQE, while the bond CF parameters remained unchanged (see Table II of SI).
 Then, the bond stretching force parameter was slightly reduced to improve the agreement of the simulated infrared (IR) absorption spectra with experimental data.
@@
 As discussed below, the incorporation of the CF term and the subsequent adjustments enable the use of a consistent set of intramolecular parameters across gas and liquid phases, with a good match to the experimental data for molecular structure properties (see Table 1).
-In addition, since adQTB and PIMD simulations lead to slightly different values for the liquid water density62, 63, two distinct vdW parameter sets were developed for each of the two methods.
+In addition, since adQTB and PIMD simulations lead to slightly different values for the liquid water density65, 66, two distinct vdW parameter sets were developed for each of the two methods.
 Unless stated otherwise, the results provided here for the Q-AMOEBA (CF) model are obtained from adQTB simulations (using the adQTB parameters set).
@@
 Liquid water simulations used 4000 molecules in a cubic box with periodic boundary conditions.
-The vdW cutoff was 12 Å, electrostatic interactions were computed with the Smooth Particle Mesh Ewald (SPME) method 71, 72 using a 7 Å real-space cutoff and a 60 × 60 × 60 grid.
-A BAOAB-RESPA integrator 73 with 0.2 fs and 2 fs time steps for bonded and nonbonded interactions, respectively, was used.
+The vdW cutoff was 12 Å, electrostatic interactions were computed with the Smooth Particle Mesh Ewald (SPME) method 74, 75 using a 7 Å real-space cutoff and a 60 × 60 × 60 grid.
+A BAOAB-RESPA integrator 76 with 0.2 fs and 2 fs time steps for bonded and nonbonded interactions, respectively, was used.
 PIMD simulations employed 32 beads with the Thermostated Ring Polymer Molecular Dynamics (TRPMD) method with a mild Langevin thermostat on the centroid with a friction coefficient of ps-1.
@@
 | Exp. | 0.957 | 104.52 | 0.97 | 105.1 |
-3 Results and Discussion
 0.1 Binding energies of Water Clusters
-Table 2 shows the Binding Energies (BE) for various water clusters, comparing the Q-AMOEBA (CF) model with AMOEBA-03 and ab initio methods 74, 76, 77.
+Table 2 shows the Binding Energies (BE) for various water clusters, comparing the Q-AMOEBA (CF) model with AMOEBA-03 and ab initio methods 77, 79, 80.
 This allows us to validate the new parameters against the original AMOEBA-03 and ab initio BE.
 The inclusion of CF has reduced the Root Mean Squared Error (RMSE) to 0.37, compared to 0.81 for AMOEBA-03.
-Table IV in SI demonstrates that Q-AMOEBA (CF) reduces the RMSE by about 25% compared to AMOEBA-03 for larger water clusters, regardless of the method (PIMD or adQTB) used for including NQEs during the parametrization.
+Table IV in SI demonstrates that Q-AMOEBA (CF) reduces the RMSE by about 25% compared to AMOEBA-03 for larger water clusters, regardless of the method (PIMD or adQTB) used for including NQE during the parametrization.
 This highlights the importance of precise monomer geometry in a flexible water model for faithfully reproducing the intricate PES of water clusters.
-3.2 Liquid properties of water
+0.2 Properties of liquid water
 In bulk liquid water, the H2O molecules’ structural dynamics generate a globally balanced and symmetric environment, resulting in most molecules having an average net charge close to zero.
@@
 As a result of this improved description of polarization effects, Q-AMOEBA (CF) predicts a change in dipole moment values from 1.84 D in the gas phase to 2.77 D in the liquid phase.
-These results align closely with both ab initio MD calculations 78 and experimental observations 79, 80, which report a gas-phase dipole moment of 1.85 D and a liquid-phase near 2.9 D.
+These results align closely with both ab initio MD calculations 81 and experimental observations 82, 83, which report a gas-phase dipole moment of 1.85 D and a liquid-phase near 2.9 D.
 In contrast, AMOEBA-03 predicts a similar value for the liquid phase (2.75 D) but underestimates the gas-phase dipole (1.77 D).
@@
 Simulations not accounting for these effects often deviate from experimental results.
-Figure 1-(a) shows IR spectra for Q-AMOEBA (CF) with a deconvolution procedure applied to correct for the spectral broadening caused by the high friction coefficient used in adQTB simulations, as previously reported 62.
-The explicit inclusion of NQE generally tends to red-shift the intramolecular peak frequencies39, which is consistent with the results of Fig.
+Figure 1-(a) shows IR spectra for Q-AMOEBA (CF) with a deconvolution procedure applied to correct for the spectral broadening caused by the high friction coefficient used in adQTB simulations, as previously reported 65.
+The explicit inclusion of NQE generally tends to red-shift the intramolecular peak frequencies36, which is consistent with the results of Fig.
 1-(a), comparing Q-AMOEBA (CF) and AMOEBA-03.
-This trend is independent on the method used for the inclusion of NQE and the PIMD curves provided in SI-Fig.4 only display marginal discrepancies with respect to the adQTB results of Fig.
+This trend is independent on the method used for the inclusion of NQE and the TRPMD curves provided in SI-Fig.4 only display marginal discrepancies with respect to the adQTB results of Fig.
 1-(a).
 Using a lower gas-phase bond stretching force in our updated FF leads to a stretching peak position (at 3400 cm-1) that closely match experimental data, without affecting liquid thermodynamic properties.
-Additionally, the CF term induces a notable red shift and intensity reduction of the angle bending peak (at 1600 cm-1) , aligning with the experimental spectrum 81, 82, 83, and in agreement with previous observations for the AMOEBA+ 25 force field.
+Additionally, the CF term induces a notable red shift and intensity reduction of the angle bending peak (at 1600 cm-1) , aligning with the experimental spectrum 84, 85, 86, and in agreement with previous observations for the AMOEBA+ 25 force field.
 Thermodynamic properties such as density, heat capacity, thermal expansion, and isothermal compressibility were assessed from 249.15 K to 369.15 K at 1 atm (see Figure 1-(b)-(g)).
@@
 NQE influence water’s density maximum due to oxygen’s electronegativity, which creates polarized OH bonds and directional hydrogen bonding that stabilizes local tetrahedral structures (similar to that found in ice).
-The observed density maximum shifts for different isotopes (277.13K for H2O, 284.34 K for D2O, and 286.55 K for T2O) 84, 39 indicate an enhanced hydrogen bond network with increasing isotope mass.
+The observed density maximum shifts for different isotopes (277.13K for H2O, 284.34 K for D2O, and 286.55 K for T2O) 87, 36 indicate an enhanced hydrogen bond network with increasing isotope mass.
 Notably, the inclusion of NQE and CF brings the predicted maximum density closer to the experimental value, with Q-AMOEBA (CF) at 279.5 K compared to AMOEBA-03’s higher value of 287 K.
-The enthalpy of vaporization (Hvap) is an other characteristic quantity that is known to be strongly affected by NQE in water.
-Q-AMOEBA (CF) achieves chemical accuracy for Hvap across all temperatures, surpassing AMOEBA-03 and previous Q-AMOEBA versions 63.
-This improvement, evident in the correct slope obtained with increasing temperature, underscores NQE’s essential role.
-Notably, H was not used in parameter fitting.
-Despite this, Q-AMOEBA (CF) accurately captures thermodynamic observables strongly influenced by NQE, demonstrating its robustness even with fewer fitted bulk properties.
-This underscores the necessity of incorporating both NQE and CF to accurately describe water’s thermodynamic properties.
-While CF’s impact on the dielectric constant is subtle (Figure 1-(g)), it significantly improves the description of molecular charge distribution.
-By enhancing the accuracy of dipole moment calculations, Q-AMOEBA (CF) demonstrates superior capability in modeling water’s local geometry and polarization effects, making it a crucial advancement in understanding water’s complex behavior.
+The enthalpy of vaporization () is an other characteristic quantity that is known to be strongly affected by NQE in water.
+Q-AMOEBA (CF) achieves chemical accuracy for across all temperatures, with a slight temperature-independent shift of kcal.mol-1, thereby improving over the previous Q-AMOEBA model (2022) 66.
+Though the absolute error of AMOEBA-03 (in classical MD simulations) is lower than that of Q-AMOEBA (CF), this is likely due to a compensation of errors since AMOEBA-03 fails to capture the temperature-dependence of the enthalpy of vaporization and exhibits a strongly overestimated slope.
+Indeed, it was shown that explicit inclusion of NQE was essential to capture the physical processes at play and recover the correct temperature slope66.
+Notably, the enthalpy of vaporization was not used as a fitting property in the parametrization, yet the combined inclusion of CF and NQE reduces the error to within 1 kcal.mol-1, while reliably reproducing the experimental temperature dependence of this NQE-sensitive thermodynamic property.
+The impact of the CF term on the estimation of the dielectric constant is shown in Figure 1-(g).
+As discussed above, CF contributes to a more accurate description of the molecular charge distribution, leading to a more accurate dipole moment calculation and a better treatment of polarization effects as a function of molecular geometry.
+While Q-AMOEBA (CF) yields a slight improvement over the original Q-AMOEBA FF in estimating the dielectric constant, a noticeable discrepancy with the experimental value remains, indicating that further refinement may be necessary.
+particular property.
 Finally, Q-AMOEBA (CF) predicts a self-diffusion coefficient of cm2/s at 298.15 K, in good agreement with experiments.
 In comparison, AMOEBA-03’s value of 2.02 × 105 cm2/s demonstrates that including NQE accelerates diffusion compared to classical dynamics.
-Note that the diffusion coefficient was calculated for Q-AMOEBA (CF) using thermostatted ring-polymer MD (TRPMD) as the adQTB method was shown in previous studies62, 63 to underestimate diffusion due to the high friction coefficients required to control zero-point energy leakage (the adQTB value for the coefficient is indeed 0.80 ± 0.01 × 105 cm2/s).
-Further studies on ice I and heavy water provided in SI expand on the role of NQE with Q-AMOEBA (CF).
-. Corresponding radial distribution functions are available in SI.
+Note that the diffusion coefficient was calculated for Q-AMOEBA (CF) using thermostated ring-polymer MD (TRPMD) as the adQTB method was shown in previous studies65, 66 to underestimate diffusion due to the high friction coefficients required to control zero-point energy leakage (the adQTB value for the coefficient is indeed 0.80 ± 0.01 × 105 cm2/s).
+Comparisons with Q-AMOEBA (2024), AMOEBA+, and AMOEBA-03 are included in SI to isolate the effects of NQE and CF individually.
+Additional studies on ice I and heavy water expand on the role of NQE with Q-AMOEBA (CF) is also available.
 0.3 Hydration Free Energy
@@
 Each thermodynamical state was thermalized for 1.5 ns and run for 5 ns.
-The Bennett Acceptance ratio (BAR) method 91 and the Path Integral BAR (PI-BAR) for PIMD results 70 was then used to estimate the free energy.
-The final free energy was taken as the sum over all windows, where 1.9 kcal.mol-1 was added for each ions 90.
+The Bennett Acceptance ratio (BAR) method 94 and the Path Integral BAR (PI-BAR) for PIMD results 73 was then used to estimate the free energy.
+The final free energy was taken as the sum over all windows, where 1.9 kcal.mol-1 was added for each ions 93.
 For water, the Q-AMOEBA (CF) model with explicit inclusion of NQE yields results that are very close to that of the original AMOEBA-03 (using classical MD), with a slight underestimation of the corresponding HFE with respect to experiment.
 Table 3, shows a clear overall improvement in the HFE values for the different ions.
-Note that for monovalent ions, this improvement was obtained with the original AMOEBA parameters90, while the Thole parameters have been slightly adjusted for Mg2+ and Ca2+ to correctly account for NQE with Q-AMOEBA (CF).
+Note that for monovalent ions, this improvement was obtained with the original AMOEBA parameters93, while the Thole parameters have been slightly adjusted for Mg2+ and Ca2+ to correctly account for NQE with Q-AMOEBA (CF).
 The new parameters are available in SI.
-To further assess the transferability of Q-AMOEBA (CF), we also proceed to compute the HFE of approximately 40 small organic compounds which were parametrized using Poltype2 92.
+To further assess the transferability of Q-AMOEBA (CF), we also proceed to compute the HFE of approximately 40 small organic compounds which were parametrized using Poltype2 95 (All parameter used for the solutes are available in the SI).
 Figure 2 displays the scatter plot of Q-AMOEBA (CF) models with AMOEBA-03 against experimental values alongside their respective RMSE.
@@
 Considering the explicit inclusion of NQE in Q-AMOEBA (CF), the higher RMSE values are reasonable, as these effects introduce additional complexity to the model.
+Furthermore, none of the solute parameters derived using PolType2 were parameterized with explicit inclusion of NQE.
+Therefore, the lower RMSE observed with AMOEBA-03 is likely due to a compensation of errors.
+Further work, focused on the role (and specific parametrization?) of the van der Waals contribution will allow to deepen our understanding of the role of NQE in hydration free energies.
 0.4 The Alanine Dipeptide
-A longstanding question in the field of biomolecular simulation concerns the quantitative and qualitative impacts of NQE 93 in proteins and in protein-ligand systems.
+A longstanding question in the field of biomolecular simulation concerns the quantitative and qualitative impacts of NQE 96 in proteins and in protein-ligand systems.
 Such calculations present inherent challenges due to the high number of conformational states accessible.
@@
 Indeed, it embodies sampling difficulties typical of larger proteins and its well known metastable states characterized by the two dihedral angles and (shown in Figure 3) serves as an ideal model for exploring NQE impacts.
-To overcome these, we first resort to the Adaptive Biasing force (ABF) technique 94, 95 using the Colvars library 96 with both the and as collective variables in order to recover the 2 dimensional free energy surface as a function of (,) with NQE using adQTB and without NQE.
+To overcome these, we first resort to the Adaptive Biasing force (ABF) technique 97, 98 using the Colvars library 99 with both the and as collective variables in order to recover the 2 dimensional free energy surface as a function of (,) with NQE using adQTB and without NQE.
 Details of the simulation, and input files are provided in SI.
@@
 | AMOEBA-03 | -12.25 |
-4 Conclusion
 In this work, we introduced Q-AMOEBA (CF), an enhanced version of the Q-AMOEBA model that incorporates a geometry-dependent CF term and explicitly accounts for NQE via the adQTB method.
@@
 It is also important to note that the vdW interactions in biological systems may require further attention in order to explicitly use NQE at large scale.
-This parametrization task will be the subject of further research in link with the biological importance of NQE in biology and will be effectively addressed using machine learning techniques, which hold great potential in training models that adapt the vdW parameters for biological contexts97.
+This parametrization task will be the subject of further research in link with the biological importance of NQE in biology and will be effectively addressed using machine learning techniques, which hold great potential in training models that adapt the vdW parameters for biological contexts100.
+Supporting Information
+All simulation parameters used in this study, including those for water, ions, and alanine dipeptide, are available in the Supporting Information (SI).
+The SI also provides the full set of Q-AMOEBA (CF) parameters, along with all input files required for reproducibility and additional figures supporting the results discussed in the main text.
+Results obtained using TRPMD simulations with Q-AMOEBA (CF), as well as the corresponding water parameters, are provided.
+In addition, the SI contains the parameters for the Q-AMOEBA (2024) water model (adQTB and TRPMD) together with figures illustrating its performance.
+Comparative results between Q-AMOEBA (CF), Q-AMOEBA (2024) (TRPMD and adQTB), and the original AMOEBA-03 model are also included.
+A detailed analysis of ice Ic and heavy water is available.
+The window parameters used for the hydration free energy calculations, as well as the newly derived Thole parameters for Ca2+ and Mg2+, are provided.
+Parameters used for all solutes studied are bundled in a tar archive.
+Finally, for ease of reproducibility, the Colvar input files used for alanine dipeptide simulations are provided, along with the PMF differences obtained with Q-AMOEBA (CF), Q-AMOEBA (2024) (adQTB and PIMD), and AMOEBA-03, as well as the decomposition of the computed solvation free energy for all models.
 Acknowledgments